At a party: exactly 15 people drank some beer, exactly 12 people drank some wine, exactly 10 people drank both beer and wine, and exactly 18 people drank neither beer nor wine. How many people at the party?

Answer by stanbon(75887)

At a party: exactly 15 people drank some beer, exactly 12 people drank some wine, exactly 10 people drank both beer and wine, and exactly 18 people drank neither beer nor wine. How many people at the party?
----
Draw 2 intersecting circles in a rectangle (Venn Diagram).
---
Put 18 in the rectangle but not in any circle.
Put 10 in the intersection of the 2 circles.
Put 15-10 = 5 in the beer circle but not in the intersection
Put 12-10 = 2 in the wine circle but not in the intersection.
---
Add up the numbers:
2 + 5 + 10 + 18 = 35 people at the party
====================
Cheers,
Stan H.
==================